OpenStudy (anonymous):

A cruise ship has circular windows in It's state rooms. What is the circumference of one of these windows If it's diameter is 15 inches ???

OpenStudy (goldphenoix):

Do you know the formula for finding the circumference of a circle?

OpenStudy (mathstudent55):

\(C = \pi d\)

OpenStudy (primeralph):



OpenStudy (goldphenoix):

Great. Now plug that number in the equation, and that would give you the ~answer~

OpenStudy (anonymous):

47.1 ?

OpenStudy (goldphenoix):

Yes. Good job. :) 47.1 inches. :)

OpenStudy (anonymous):

I tried that it was wrong

OpenStudy (anonymous):

@GoldPhenoix

OpenStudy (goldphenoix):

That should be the answer. Does it say anything about rounding?

OpenStudy (anonymous):

no it doesn't. Your not supposed to round unless the problem says too and it doesn't tell you too

OpenStudy (mathstudent55):

\(C = 15 \pi \)

OpenStudy (goldphenoix):

^That can work, since that is the exact value.

OpenStudy (anonymous):

i don't have the pi sign on my computer

OpenStudy (mathstudent55):

Leave \(\pi\) in the answer to have the exact answer.

OpenStudy (anonymous):

and it wont let me copy and paste

OpenStudy (mathstudent55):

Write pi

OpenStudy (mathstudent55):

15pi or 15(pi)

OpenStudy (anonymous):

ok let mee see if it works

OpenStudy (mathstudent55):

Also, does it need units? 15pi inches or 47.1 inches

OpenStudy (anonymous):

all wrong

OpenStudy (anonymous):

?

OpenStudy (goldphenoix):

Hmm. Are you sure you write the question correctly?

OpenStudy (anonymous):

yess!!!

OpenStudy (anonymous):

positive